OpenStudy (anonymous):

Identify the conjugate of 26-6√14 and explain, in complete sentences, why it is the conjugate.

OpenStudy (anonymous):

conjugate of \[a+\sqrt{b}\] is \[a-\sqrt{b}\] so your conjugate is \[26+6\sqrt{14}\]

OpenStudy (anonymous):

yeah i got that but it tells me to explain why that is its conjugate